the base of a triangle exceeds the height by 9 feet. if the area is 315 square feet, find the length of the base and the height of the triangle

Answer by nerdybill(7384)

the base of a triangle exceeds the height by 9 feet. if the area is 315 square feet, find the length of the base and the height of the triangle
.
Let h = height
then
h+9 = base
.
since area of triangle = (1/2)bh
we have
315 = (1/2)(h+9)(h)
630 = (h+9)(h)
630 = h^2+9h
0 = h^2+9h-630
factoring:
0 = (h+30)(h-21)
h = {-30, 21}
throw out the negative solution (extraneous) leaving:
h = 21 feet (height)
.
base:
h+9 = 21+9 = 30 feet